Order VPS Hosting
Order a VPS, Semi- dedicated or Dedicated server in Dallas, London or Australia.

Get Assistance
Ask our support team about your hosting requirements.


Host where the staff takes pride in making customers happy

I don't know what to say.... I'm a bit stunned, this is not the kind of thing that happens to me. I really appreciate it and I'll be recommending you.

- Jerry (After we worked out some custom pricing for him. If you want 'stunning' customer service too, then become a RimuHosting server owner.) (#190/332)
Home > Support > Notices

Related Links

Notice Links:

Notices

London switch restart

PostedFri, 18 Jun 2021 05:00 AM UTC
Last UpdateWed, 23 Jun 2021 20:38 PM UTC (33 days ago)
StatusOpen

Wed, 23 Jun 2021 20:35 PM UTC: due to recuring issues with the old switch, under urgency connections have been moved to the new switch and all services restored there. Systems look stable presently, we will continue to monitor.

Wed, 23 Jun 2021 02:01 AM UTC: this switch restarted itself again. It has a possibly failing power supply. We have a replacement switch in the rack and we are actively working on preparing that so we can move affected services to that.

One of our cabinet switches in London restarted itself. This impacted service to 7 connected hosts for 2-3 minutes. The switch is back online and is passing traffic normally again.

We are continuing to monitor this switch and will make further checks on the cause of the outage.

#

Dallas UPS battery maintenance

PostedWed, 9 Jun 2021 00:53 AM UTC
Last UpdateWed, 9 Jun 2021 00:53 AM UTC (7 weeks ago)
StatusOpen

The datacenter team at our Dallas site are performing battery replacement in selected UPS units as a followup to previous inspections.

No impact is expected

Maintenance Window: 2021-06-21 08:00:00 EDT - 2021-06-23 04:00:00 EDT

#

London network packet loss

PostedMon, 28 Jun 2021 16:33 PM UTC
Last UpdateTue, 29 Jun 2021 03:31 AM UTC (28 days ago)
StatusClosed
Affected Data CenterLondon

Mon, 28 Jun 2021 21:36 PM UTC: No further issues have been seen. We are documenting a mitigation plan for future events like this.

Mon, 28 Jun 2021 17:22 PM UTC: The traffic appear to have mostly subsided now, and normal service is restored. We are continuing to monitor this and working further with our upstream.

Mon, 28 Jun 2021 16:53 PM UTC: There is a traffic flood targeting one one of our customers. We are working to identify the target and isolate that from our network to clear things up for everyone else while we investigate further.

We are seeing intermittent connectivity issues affecting our London site. Working with our upstream to discover and resolve the issue.

#